40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 24 00:39:47 2017, gautam, Update, LSC, DRMI locking attempt 
    Reply  Fri Aug 25 01:20:52 2017, gautam, Update, LSC, DRMI locking attempt 
       Reply  Wed Aug 30 19:49:33 2017, gautam, Update, LSC, REFL55 demod board debugging REFL55_demod_check.pdf
          Reply  Thu Aug 31 00:52:52 2017, gautam, Update, LSC, REFL55 whitening board debugging REFL55_whtCheck.pdfREFL55_whtChk.tar.gz
             Reply  Thu Aug 31 03:31:15 2017, gautam, Update, LSC, DRMI re-locked! DRMI_1f.pngDRMI_relocked.pdf
                Reply  Mon Sep 4 16:30:06 2017, gautam, Update, LSC, Oplev loop tweaking 
                   Reply  Mon Sep 4 18:18:29 2017, rana, Update, LSC, dewhite switching: FOTON settings 
                   Reply  Tue Sep 5 02:07:49 2017, gautam, Update, LSC, Low Noise DRMI attempt MICH_err_comp.pdfdeWhitenedCoil.pdf
                      Reply  Wed Sep 6 23:06:30 2017, gautam, Update, LSC, Coil de-whitening switching investigation BScoils.pdfITMcoils.pdf
                         Reply  Fri Sep 15 16:00:33 2017, gautam, Update, LSC, Sensing measurement DRMI1f_Sep5.pdf
                         Reply  Fri Sep 15 17:08:58 2017, gautam, Update, LSC, Coil de-whitening switching investigation DRMI_IntensityNoise.pdf
                            Reply  Sat Sep 16 10:56:19 2017, rana, Update, LSC, Coil de-whitening switching investigation 
                               Reply  Fri Sep 22 18:12:27 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                  Reply  Wed Sep 27 22:25:21 2017, gautam, Update, LSC, DAC noise measurement (again) coilNoises.pdf
                                     Reply  Thu Sep 28 06:35:44 2017, Chris, Update, LSC, DAC noise measurement (again) 
                                        Reply  Thu Sep 28 23:50:04 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                           Reply  Fri Sep 29 18:36:25 2017, gautam, Update, LSC, DAC noise measurement (again) 
                               Reply  Tue Oct 3 01:32:39 2017, gautam, Update, LSC, Laser intensity noise coupling to MICH (simulated) MICH_intensityNoiseCoupling.pdfMICH_intensityNoiseASD.pdf
Message ID: 13314     Entry time: Fri Sep 15 17:08:58 2017     In reply to: 13300     Reply to this: 13315
Author: gautam 
Type: Update 
Category: LSC 
Subject: Coil de-whitening switching investigation 

I downloaded a segment of data from the time when the DRMI was locked with the BS and ITM coil driver de-whitening switched on, and looked at coherence between MC transmission and the MICH error signal. Attachment #1 doesn't show any broadband high coherence between 60-300Hz, so it cannot explain the noise in the full range between 60-300Hz. 

The DQ channel for the MC transmission is recorded at 1024 kHz, so to calculate the coherence, I had to decimate the 16K MICH data. 

Since we have the AOM installed, I suppose we can actually measure the intensity noise coupling to MICH by driving a line in the AOM. 

I also checked for coherence in the 60-300Hz band between MICH/PRCL and MICH/SRCL, and didn't see any appreciable coherence. Need to think about this more.

Quote:

 Rana suggested checking coherence with MC transmission to see if this could be laser intensity noise.

Attachment 1: DRMI_IntensityNoise.pdf  196 kB  | Hide | Hide all
DRMI_IntensityNoise.pdf
ELOG V3.1.3-